Ebenezer Scrooge has invested 60% of his money in share A and the remainder in share B. He assesses their prospects as follows:

A B
Expected return (%) 15 20
Standard deviation (%) 20 22
Correlation between returns .5

a. What are the expected return and standard deviation of returns on his portfolio? (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)

Expected return 17.00 correct %
Standard deviation 18.08 correct %

b. How would your answer change if the correlation coefficient were 0 or –.5?
Answer #1

Wa= Weight of investment in stock A 60% OR 0.60

Wb= Weight of investment in stock A 40% OR 0.40

Ra = Return of Stock A

Rb = Return of Stock B

SDa = Standard Deviation of Stock A

SDb is the Standard Deviation of Stock B

R is the correlation between Stock A and B

Expected Return = Ra*Wa + Rb*Wb

= 15% * 0.60 + 20 * 0.40

= 17%

SD of Portfolio = \large \sqrt{(SDa)^2*Wa^2+(SDb)^2*Wb^2 + 2*R*SDa*Wa*SDb*Wb}

=  \bg_white \large \sqrt{(.20)^2*0.60^2+(0.22)^2*0.40^2 + 2*0.5*.20*0.60*0.22*0.40}

= 0.1808 OR 18.08%

SD of Portfolio When R = 0

SD of Portfolio = \large \sqrt{(SDa)^2*Wa^2+(SDb)^2*Wb^2 + 2*R*SDa*Wa*SDb*Wb}

=\bg_white \large \sqrt{(SDa)^2*Wa^2+(SDb)^2*Wb^2 + 2*0*SDa*Wa*SDb*Wb}

= \bg_white \large \sqrt{(SDa)^2*Wa^2+(SDb)^2*Wb^2 + 0

=  \bg_white \large \sqrt{(.20)^2*0.60^2+(0.22)^2*0.40^2

= 0.02214 OR 2.214%

SD when R= -5

SD of Portfolio = \large \sqrt{(SDa)^2*Wa^2+(SDb)^2*Wb^2 + 2*R*SDa*Wa*SDb*Wb}

=  \bg_white \large \sqrt{(.20)^2*0.60^2+(0.22)^2*0.40^2 + 2*(-0.5)*.20*0.60*0.22*0.40}

= \bg_white \large \sqrt{0.0144+ 0.007744- 0.01056}

= \bg_white \large \sqrt{0.011594}

= 0.1076 OR 10.76%

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.